Determine the open intervals on which the graph of f(x)= (x²)/(x²+1) is concave upward or concave downward


1
Expert's answer
2021-12-02T16:57:38-0500
"f(x)=\dfrac{x^2}{x^2+1}"

Domain: "(-\infin, \infin)"


"f'(x)=\dfrac{2x(x^2+1)-2x(x^2)}{(x^2+1)^2}=\dfrac{2x}{(x^2+1)^2}"

"f''(x)=\dfrac{2((x^2+1)^2-2x(2x)(x^2+1))}{(x^2+1)^4}"

"=\dfrac{2(x^2+1-4x^2)}{(x^2+1)^3}=\dfrac{2(1-3x^2)}{(x^2+1)^3}"

Find the point(s) of inflection


"f''(x)=0=>\dfrac{2(1-3x^2)}{(x^2+1)^3}=0"

"1-3x^2=0"

"x_1=-\dfrac{\sqrt{3}}{3}, x_2=\dfrac{\sqrt{3}}{3}"

If "x<-\dfrac{\sqrt{3}}{3}, f''(x)<0, f(x)" is concave downward.

If "-\dfrac{\sqrt{3}}{3}<x<\dfrac{\sqrt{3}}{3}, f''(x)>0, f(x)" is concave upward.

If "x>\dfrac{\sqrt{3}}{3}, f''(x)<0, f(x)" is concave downward.


"f(x)" is concave upward on "(-\dfrac{\sqrt{3}}{3}, \dfrac{\sqrt{3}}{3})."

"f(x)" is concave downward on "(-\infin,-\dfrac{\sqrt{3}}{3})\cup ( \dfrac{\sqrt{3}}{3}, \infin)."